# Interactive Program Synthesis for Modeling Collaborative Physical Activities from Narrated Demonstrations

### v1

@@ 6.4. Initial System Learning from Narrated Demonstrations of Users (RQ2)
Pre-feedback alignment of learned programs is shown in Fig.7 (top left) and Fig.7 (bottom left). The wide variance likely reflects differences in instructional quality. Among the 10 participants whose programs achieved correctness or completeness scores greater than 80%, their narrations shared two traits: (i) 9/10 explicitly stated the purpose behind actions (e.g., Now your teammate will make an overlapping run, and you want to pass it to them so that they can score the goal.” (P5)), and (ii) 7/10 articulated multiple possible outcomes to keep executions correct under varied conditions (e.g., We’re gonna see if the shot is open, I’m gonna take it, but if it’s not, now I’m gonna pass it. Uh, or, let’s say I got covered, I could dribble a bit, and then I can take the shot.” (P15)). In contrast, 7 participants whose learned program scored less than 50% in either metrics exhibited the opposite, either did not explain the rationale (3/7) or did not consider multiple outcomes (4/7).
@@ 6.5. System Learning from Corrective Feedback of Users (RQ3)
6 of 20 participants did not improve on either correctness or completeness after feedback. These outcomes stemmed from three sources: (1) Insufficient APIs: 2 out of 7 participants tried to correct the trajectory curvature of movements taken by the program-controlled avatar. However, our APIs supported only optimal path planning/control and, therefore, could not accommodate this feedback. (2) Context loss during repair: To reduce synthesis time, we omitted original narrated videos from the LLM repair prompt for feedback on system’s narrated demonstrations. Two of 20 participants then gave feedback that seemingly contradicted their earlier teaching (assuming the system retained that context), leading the LLM to delete intended branches. For example, a unique issue occurred for P17, who ended feedback with “Okay, we can stop here,” which the LLM took literally, producing a truncated program. This accounted for 2/7 non-improving cases. (3) Deadlocks: In 4 of 7 participants, the synthesized program deadlocked because a learned condition was never satisfied. Our system did not explain which condition blocked execution, preventing effective debugging by the users.
@@ 7. Discussion
While the flexibility of free-form narrated demonstrations could allow users to teach their unique strategies, we observed that this freedom could also introduce challenges in externalizing tacit physical knowledge. The initial program synthesis exhibited wide variability, with standard deviations of 24.02% for correctness and 28.85% for completeness. Our analysis suggests this variability stems from differences in teaching styles; participants who explicitly articulated the rationale behind their actions and described potential contingencies tended to produce programs that were more aligned with their intent. Conversely, programs synthesized from narrated demonstrations that omitted this contextual information were tend to perform poorly.
In future work, this omitted information could be elicited by incorporating a framework inspired by Theory of Mind (Premack and Woodruff, 1978; Baker et al., 2017; Rabinowitz et al., 2018). Instead of passively receiving instruction, the system could actively reason about the user’s mental model of the task. For instance, an LLM could analyze the narrated demonstrations to identify points of ambiguity or underspecified conditions. It could then generate targeted questions to probe the user’s intent, such as, “I noticed you moved left. Was that to create space, or were you reacting to your teammate’s position?” By modeling the user’s knowledge and identifying gaps, the system could guide the user to provide the contingencies and rationales needed for robust program synthesis, shifting the interaction from a one-way demonstration to a collaborative dialogue.
Our results demonstrate that participants were able to successfully interpret and correct the synthesized programs to better align them with their intent. This corrective feedback loop proved effective; even with a single round of feedback—which 16 of 20 participants provided due to time constraints—program completeness increased significantly by 16.68% (SD=18.52, p=0.03). This quantitative improvement was echoed in participants’ subjective ratings of the decision-flow structure, which also improved post-feedback (Fig. 8). Participants were consistently able to steer the synthesized program toward their intended high-level logic for the soccer tactics (Fig. 7, right). In interviews, several participants noted that additional, iterative rounds of feedback—rather than a single round—would likely yield further improvements.
However, the study also revealed key limitations. Unlike completeness outcomes, the improvement in program correctness showed upward trend of 6.87% (SD=22.42) on average but was not statistically significant (p=0.10). Our analysis pointed to two fundamental issues. First, the expressiveness of the underlying API library was a significant constraint. In some cases, participants were unable to improve program alignment because the provided APIs lacked the necessary functions to implement their intended corrections. Recall that our system restricted LVLM to not arbitrarily implement APIs and only use the ones provided. While this restriction may be helpful in reducing potential hallucinations of LVLM, it could also constrain the expressivity in modeling user behavior in trade-off. This suggests that for a system to be broadly applicable within a domain, its API library must be iteratively developed through an in-depth analysis of user needs.
Another major limitation of the system was its failure to explain program deadlocks. On 4 out of 6 occasions where users fail to better align the programs with their intent, users experienced deadlocks where the programs halted in the middle because one of its preconditions are not satisfied. The system could only narrate the program logic and lacked a mechanism to detect and communicate which specific condition(s) were causing the halt. This made it particularly challenging for users to debug when preconditions involve multiple constraints. This limitation could be mitigated by augmenting the synthesis prompt to generate routines specifically for detecting deadlocks and communicating the source of the error to the user.
While the majority of participants (95%) found the narrated demonstration interface easy to use, their feedback also highlighted a key question regarding the necessary scope of instruction for collaborative tasks. Our study utilized environment agents with pre-scripted behaviors, which several participants found constraining when the agents’ actions diverged from their tactical intent. This suggests that for complex collaborative activities, modeling the user’s behavior in isolation may be insufficient for fully externalizing their knowledge. A natural extension of our work would be to allow users to teach or modify the behavior of other agents in the environment. While our system’s architecture could be adapted for this purpose, such an extension would introduce additional system complexity and potentially increase the user’s cognitive and physical load, requiring careful investigation into mitigation strategies.
While the majority of participants (90%; 18/20) found the process of correcting the model to be straightforward, a key challenge emerged in understanding what the model had learned, which in turn complicated debugging. Five participants reported that the decision-flow diagrams became difficult to interpret as they grew in complexity, often containing redundant nodes and edges. This issue likely stems from the absence of a parsimony objective during program synthesis and visualization. Future iterations could improve interpretability by imposing minimality constraints on the program structure and its visual representation within the LLM prompt. Furthermore, while most participants inferred behavior from the system-generated narrated demonstrations, some were confused by technical jargon (e.g., “check for a path width of 2 meters”). This highlights the need to extend the interface to allow users to inquire about and clarify ambiguous terms, making the system’s communication more transparent.
8. Conclusion
In this paper, we presented an interactive program synthesis system that enables non-programmers to teach complex, collaborative physical tasks. We framed this problem as one of program synthesis, where the system learns an editable, interpretable program that represents the collaborative activity. Our core contribution is an interaction model that uses narrated demonstrations as a unified, bidirectional modality: users teach and correct the system through paired physical actions and language, and the system communicates its learning using the same modality, removing the need for users to interact with raw code. Our study showed feasibility of this approach. Users successfully taught multi-agent soccer tactics and significantly refined the synthesized programs to align with their intent. Our findings have implications for the design of future interactive systems for task learning, highlighting the need for mechanisms to more actively elicit tacit knowledge from users and to expand the instructional scope to include the behavior of collaborating agents. This work contributes a novel interaction framework and demonstrates the feasibility of synthesizing executable models of complex human collaboration directly from end-user instruction.
@@ A.5. Generalizability of Learned Programs
The learned Scenic probabilistic programs did not simply replay participants’ narrated demonstrations. In the study, each participant provided two demonstrations of the same task, and these variations were modeled as distributions in Scenic programs. Semantically, a Scenic execution samples a concrete scenario from the distribution it encodes and generates it in simulation. Fig. 9 overlays two demonstration trajectories of a participant with five program execution trajectories for each of the three scenarios: overlap (left), lure (middle), and distribute (right), corresponding to Participants 9, 10, and 11. These participants produced learned programs that achieved 100% correctness and completeness via interactive program synthesis.
A.6. System Implementation
A.6.1. Teaching, Inspection, Editing Interfaces
@@
A.6.2. APIs
Our study personnels consisted of soccer players with the history of playing in the competitive soccer leagues. We used our domain knowledge of soccer to model the sets of action space and the physical constraints as APIs. For action space, we provided APIs that execute moving/dribbling, passing, shooting a goal, and triggering a teammate to pass the ball to the user. For modeling the physical constraints, we APIs that can validate conditions such as (1) ball possession, (2) distance relations, (3) horizontal/vertical relations, and (4) existence of passing lane between two players of interest. To validate these conditions, the APIs were internally given access to the world states (e.g. positions, orientations, ball possession of each player) in mixed reality simulation at each timestep. The annotations of all APIs were provided to the large vision language model (LVLM) for program synthesis. For example, Fig. 10 shows the annotation of DistanceTo() API which checks for the distance relation between two players of interest.
A.6.3. System Prompt
@@
Appendix B Applicability of Our System to Other Collaborative Physical Domains
Our system architecture could support other collaborative physical domains where agents must coordinate in space and time, because it decouples domain-agnostic synthesis/inspection/repair loops from domain-specific action and constraint primitives. Narrated demonstrations are still used in the teaching interface; the LVLM compiles multimodal traces into Scenic programs whose stochastic scene specifications capture variation in strategies and environments. Users then inspect and correct behaviors via decision-flow visualization and narrated executions. Retargeting to another domain entails (i) modeling the target environment and agents in Scenic and (ii) replacing the API library.
Example in Manufacturing Domain
We keep the synthesis, inspection, and repair loop unchanged and swap the API library for this domain with three actions and four constraints. In this scenario, the user provides teaching for a humanoid robot as to how to collaborate with a human worker.
Action APIs: goTo(location), pick(item, from),
swapBuckets(new_bucket, old_bucket, station).
Constraint APIs:
BelowThreshold(bucket, n):
Trigger predicate from vision/weight sensing; true when the measured part count in bucket is <n<n.
HumanReady(worker):
returns true when the worker signals it is safe to perform the swap.
Instruction from the narrated demo (“when the human worker’s bucket containing assembly parts has less than 10 parts, go fetch a new bucket full of the assembly parts from the supply station and swap the human worker’s bucket under permission”):
behavior UserBehavior():
while True:
  if BelowThreshold(worker_bucket, 10):
    do Speak(‘‘the worker’s bucket is running low on assembly parts. Fetch another bucket from the supply station’’)
do goTo(SupplyStation)
do Speak(‘‘pick up another bucket full of parts’’)
do pick(full_bucket, SupplyStation)
    do Speak(‘‘Return to the worker’s station.’’)
do goTo(WorkerStation)
    do Speak(‘‘Wait until the worker permits the bucket swap.’’)
do Wait() until HumanReady(worker)
    do Speak(‘‘The worker gave permission. Swap the bucket.’’)
do swapBuckets(full_bucket, worker_bucket, WorkerStation)

### v2

@@ 7.4. Initial System Learning from Narrated Demonstrations of Users (RQ2)
Pre-feedback alignment of learned programs is shown in Fig.9 (first plot from the left) and Fig.9 (third plot from the left). The wide variance likely reflects differences in instructional quality. Among the 10 participants whose programs achieved correctness or completeness scores greater than 80%, their narrations shared two traits: (i) 9/10 explicitly stated the purpose behind actions (e.g., Now your teammate will make an overlapping run, and you want to pass it to them so that they can score the goal.” (P5)), and (ii) 7/10 articulated multiple possible outcomes to keep executions correct under varied conditions (e.g., We’re gonna see if the shot is open, I’m gonna take it, but if it’s not, now I’m gonna pass it. Uh, or, let’s say I got covered, I could dribble a bit, and then I can take the shot.” (P15)). In contrast, 7 participants whose learned program scored less than 50% in either metrics exhibited the opposite, either did not explain the rationale (3/7) or did not consider multiple outcomes (4/7).
@@ 7.5. System Learning from Corrective Feedback of Users (RQ3)
7 of 20 participants did not improve on either correctness or completeness after feedback. These outcomes stemmed from three sources: (1) Insufficient APIs: 2 out of 7 participants tried to correct the trajectory curvature of movements taken by the program-controlled avatar. However, our APIs supported only optimal path planning/control and, therefore, could not accommodate this feedback. (2) Context loss during repair: To reduce synthesis time, we omitted original narrated videos from the LLM repair prompt for feedback on system’s narrated demonstrations. 2 out of 7 participants then gave feedback that seemingly contradicted their earlier teaching (assuming the system retained that context), leading the LLM to delete intended branches. For example, a unique issue occurred for P17, who ended feedback with “Okay, we can stop here,” which the LLM took literally, producing a truncated program. This accounted for 2/7 non-improving cases. (3) Deadlocks: In 4 of 7 participants, the synthesized program deadlocked because a learned condition was never satisfied. Our system did not explain which condition blocked execution, preventing effective debugging by the users.
@@ 8. Discussion
In this section, we present our analysis of the user study and discuss the limitations of our system.
8.1. Learning Collaborative Physical Activities Needs Iterative Corrections
@@
We observed frequent misinterpretations of users’ intent. A recurring pattern was that the vision language model (VLM) modeled narrated content reasonably well but struggled to infer intent from visual demonstrations across the three tactical scenarios. For example, in the lure scenario (also used as the running example in Sec. 4), a user moved to the side of the field while remaining near a teammate with the ball and narrated, “to lure the opponent to the side of the field, move around here.” The narration specified moving to the side, and the demonstration implicitly indicated how far from the teammate to position. The intent was to stay close to the teammate. However, the VLM captured only the high-level instruction “move to the side of the field,” allowing positions anywhere on that side, including locations near the goal far from the teammate.
We also observed variability in users’ teaching quality, which contributed to the substantial variability in the accuracy of the initially generated programs (SD = 24.02% for correctness; 28.85% for completeness). Our analysis indicates that participants who articulated the rationale behind their actions and specified contingencies with respect to surrounding agents’ behaviors produced system learning that better matched their intent. In contrast, users who merely instructed a sequence of actions while omitting such contextual information yielded less accurate models. Furthermore, in a few instances, we observed users unintentionally contradict their initial teaching during the feedback phase, which further degraded system learning. Differences in participants’ coaching experience (Table 2) may partially explain this variability in teaching quality.
8.2. On the Effectiveness of the Interfaces to Teach, Inspect, and Correct System Learning
Given the limitations of both the generative model and users’ teaching, we now examine how our interfaces help users teach, inspect, and correct the system’s learning. Our system architecture provides a unified interaction modality: narrated demonstrations serve as the modality for teaching, for the system’s explanations of its learning, and for users’ corrective feedback. Participants generally perceived this modality as natural and easy to use. As shown in Fig. 8, 95% (19/20) of users reported that it was easy to teach collaborative activities to our system. 80% (16/20) found the decision flow easy to understand, and all participants reached a consensus that the system’s narrated demonstrations were easy to understand. Finally, 90% (18/20) indicated that it was easy to correct the system’s learning.
Beyond perceived usability, users were able to leverage these interfaces to correct the system’s learning and help it converge toward more accurate models of collaborative physical activities. Even with only a single round of feedback for most participants (16/20, due to time constraints), program completeness increased by 16.68% on average (SD = 18.52, $p=0.03$). Correctness also showed a positive, though non-significant, gain of 6.87% (SD = 22.42, $p=0.10$). These quantitative improvements were mirrored in participants’ subjective ratings of their ability to steer the modeled programs toward the intended soccer tactics (Fig. 10). In interviews, several participants further noted that additional iterative rounds of feedback would likely have yielded even more accurate program representations.
8.3. Limitations
Our approach currently lacks mechanisms for detecting inconsistencies, ambiguities, or under-specification in users’ instructions to better elicit their domain knowledge. The system takes narrated demonstrations at face value when generating or repairing programs. As a result, the system amplified individual differences in teaching ability, leading to wide variation in the accuracy of its learned models. Furthermore, the observed issues with user inspection raise questions about the scalability of our system. Twenty-five percent (5/20) of participants reported that the decision-flow diagrams became difficult to interpret as they grew in complexity, often containing redundant nodes and edges. This problem is likely to worsen as the number of agents or the complexity of activities increases. One potential workaround is to hide the decision-flow diagrams and rely solely on system executions for feedback. However, this would require users to watch too many executions to cover all branching code paths induced by contingencies, underscoring the need for mechanisms that support holistic inspection of program structure without cognitively overloading users to solicit feedback.
Although users’ feedback improved the average correctness of system learning, it also increased fluctuations in correctness outcomes after feedback, i.e., standard deviation increased (Fig. 9, left). Beyond individual variation in instruction quality, our analysis points to two major sources of error. First, the expressiveness of the API library was a major constraint. While the APIs enabled the insertion of domain knowledge to facilitate learning from only a few narrated demonstrations, they also restricted the spectrum of behaviors that could be modeled. For example, one user attempted to teach an avatar to follow a curved trajectory, which our APIs did not model; as a result, the avatar moved in a straight line instead. Second, our inspection interface did not support handling deadlocks in program execution. In 10% (2/20) of user evaluations, programs halted during user inspection phase because none of their preconditions were satisfied. The system could narrate the program logic but lacked mechanisms to detect and communicate to users that its halted because certain preconditions are not satisfied.
9. Conclusion & Future Work
Modeling collaborative physical activities has broad implications for HCI, from behavior modeling to tutorial and assistance systems. This makes the design of interaction techniques for iteratively correcting models of collaborative activities an important area for HCI research. In this work, we present an interactive system that enables users with no programming experience to author programs that model collaborative activities as intended. The physical nature of these activities requires interfaces that not only support teaching them, but also enable users to inspect and correct the system’s learning. We therefore propose interfaces that use narrated demonstrations as a unified, bidirectional modality: users teach and correct the system through paired physical demonstrations and natural language, and the system communicates its learned behavior through the same modality, without exposing raw code. Our user study in a tactical soccer setting, together with an existence proof in a factory environment, demonstrates the feasibility of this approach. Taken together, our findings point toward interactive task-learning systems that leverage interpretable program representations and human-in-the-loop refinement to capture collaborative behaviors from user instructions.
Our study surfaces key design insights that future systems should account for. It shows that programmatically embedding domain knowledge via APIs can enable learning complex collaborative activities from only a handful of narrated demonstrations, despite the difficulty of this problem. However, these APIs must be carefully designed through in-depth analysis of user needs and practices in each target domain to support a broad spectrum of users. Moreover, while modeling collaborative activities as programs makes it easy to insert domain knowledge, it can also introduce deadlocks. To help users localize and correct such situations, underlying generative models could be further instructed to create code routines for detecting deadlocks and explicitly articulating the sources of these errors to users for correction.
The study results also raise open questions for the HCI community. One concerns scalability, particularly in relation to decision-flow diagrams. Although these diagrams are useful for holistically visualizing overall program structure, they become increasingly difficult for users to comprehend as the complexity of activities or the number of agents grows. One potential direction is to display hierarchical representations, abstracting subsets of nodes and edges into higher-level units to reduce visual complexity. However, algorithms for determining appropriate levels of abstraction as a function of task complexity remain to be investigated.
Another open question relates to detecting inconsistencies, ambiguities, or incompleteness in user instructions in order to better elicit accurate domain knowledge. It may be fruitful to incorporate formal languages and algorithms from formal methods (Clarke2018HandbookModelChecking) to mathematically formalize the specifications in users’ instructions and detect logical inconsistencies, ambiguity, or under-specification. At the same time, the expressiveness of existing formal languages for capturing relevant spatial and temporal constraints needs to be examined, along with whether the underlying algorithms can be executed efficiently enough to support real-time inspection and maintain high usability. Another promising direction is to incorporate agentic reasoning, in which LLM-based agents proactively infer missing details and resolve inconsistencies or ambiguities in user instructions based on context (Plaat2025AgenticLLM).
@@ A.5. Generalizability of Learned Programs
The learned Scenic probabilistic programs did not simply replay participants’ narrated demonstrations. In the study, each participant provided two demonstrations of the same task, and these variations were modeled as distributions in Scenic programs. Semantically, a Scenic execution samples a concrete scenario from the distribution it encodes and generates it in simulation. Fig. 11 overlays two demonstration trajectories of a participant with five program execution trajectories for each of the three scenarios: overlap (left), lure (middle), and distribute (right), corresponding to Participants 9, 10, and 11. These participants produced learned programs that achieved 100% correctness and completeness via interactive program generation.
A.6. Diversity of User-Authored Programs
For each scenario, a rubric with three to four task requirements were provided (Appendix A.3), but these were potentially many ways to complete them. As a result, users authored programs reflecting diverse strategies per scenario. For example, in the lure scenario, some participants used fixed spatial thresholds (e.g., specific distances or lanes) while others encoded more conditional, temporal logic (e.g., “if the defender steps, then pass; otherwise dribble or reset”). In the overlap scenario, some prioritized getting wide first, another on dribbling towards goal immediately, and others focused on the exact shape and timing of the run, leading to different code branching structures and termination conditions. In the distribute scenario, some users first scanned the field to identify open teammate, while others moved around to create open lanes for a teammate. Together, these variations show that our technique supported users to externalize diverse strategies reflecting their own domain knowledge.
A.7. System Implementation
A.7.1. Teaching, Inspection, Editing Interfaces
@@
A.7.2. APIs
Our study personnels consisted of soccer players with the history of playing in the competitive soccer leagues. We used our domain knowledge of soccer to model the sets of action space and the physical constraints as APIs. For action space, we provided APIs that execute moving/dribbling, passing, shooting a goal, and triggering a teammate to pass the ball to the user. For modeling the physical constraints, we APIs that can validate conditions such as (1) ball possession, (2) distance relations, (3) horizontal/vertical relations, and (4) existence of passing lane between two players of interest. To validate these conditions, the APIs were internally given access to the world states (e.g. positions, orientations, ball possession of each player) in mixed reality simulation at each timestep. The annotations of all APIs were provided to the large vision language model (LVLM) for program generation. For example, Fig. 12 shows the annotation of DistanceTo() API which checks for the distance relation between two players of interest.
A.7.3. System Prompt
@@
Appendix B Existence Proof in Factory Environment
B.1. Factory Tasks and Scenarios
In the factory setting, participants taught collaborative box-packaging interactions between a mobile robot and a human worker. Unlike the soccer study, the scenarios were not pre-scripted: before teaching, each participant authored the initial scene from scratch by placing the human, robot, boxes, and other objects anywhere they wished in the Unity environment. Each participant was assigned one of two scenarios designed to exercise complementary sides of the collaboration rubric. In the robot-focused scenario (robot as learner), the participant’s demonstrations focused on teaching the robot how to collaborate with the human worker, by fetching and delivering boxes to support the human’s packaging workflow. In the human-focused scenario (human as learner), the participant’s demonstrations focused on teaching the human how to collaborate with the robot, by deciding when and how to signal the robot to bring a new box and then packaging the box once it arrived.
B.2. Study Protocol
The existence proof followed the same study procedure as the main soccer study (Appendix A), with modifications to support authoring-from-scratch in the factory setting. We recruited three participants for the factory existence proof following the same recruitment protocol as in the study 6.1, without imposing any inclusion criteria. Before teaching, we provided a short live tutorial demonstrating the new authoring-from-scratch interface: how to place agents and objects in the scene and what actions the agents could take, how to start/stop recording demonstrations, how to provide feedback to the decision flow diagram and system execution of the generated program, and how to use the rubric to evaluate behavior.
B.3. Rubrics for Factory Scenarios
As in the main study, participants used the rubric associated with their scenario to judge whether the system’s learned behavior matched their intent.
Robot-focused scenario rubric (teaching the robot).
  1. (1)
    Robot fetches the box.
  2. (2)
    Robot moves to the correct position next to the human to deliver the box.
  3. (3)
    Robot puts down the box.
  4. (4)
    Human is able to package the box.
Human-focused scenario rubric (teaching the human).
  1. (1)
    Human signals for the robot to bring a new box.
  2. (2)
    Robot brings box to human.
  3. (3)
    Human packages the new box.
As in the soccer setting, each rubric item was rated on a 0–2 scale (0 = not achieved, 1 = partially achieved, 2 = fully achieved).
B.4. System Implementation for the Factory Setting
B.4.1. Action and Constraint APIs
The factory setting extends the domain-specific API library described in the main text with actions and constraints tailored to human–robot box-packaging. Concretely, we provide action APIs such as moving to a target position, picking up an object, putting down an object, packaging a box, and raising a hand. In addition to the spatial and logical constraints used in the soccer setting (e.g., distances, regions, and compositional logic), the factory environment includes constraints that reason about object and agent state, such as whether a box is currently held or already packaged and whether a particular agent (e.g., the human) has raised their hand to request assistance. These APIs allow the system to express and infer conditions such as “the human has requested a new box”, “the robot is holding a box”, or “the box has been fully packaged”.
B.4.2. Authoring from Scratch via LLM-Generated Scenes
A key difference from the soccer setting is the authoring-from-scratch workflow on the desktop instead of demonstrating on pre-scripted scenarios in the mixed reality environment. In the factory setting, each time the participant presses the Record button to speak to author a new scene and demonstration, the system:
  1. (1)
    Sends a prompt to a language model that includes:
    • the current list of available prefabs (e.g., human, robot, box),
    • the catalog of available actions (as above),
    • any clicked positions or selected objects from the interface.
  2. (2)
    The LLM responds with a JSON specification that either:
    • spawns new objects in the scene (using prefabs and world positions), or
    • updates existing objects by assigning them actions to execute.
At this stage, the LLM’s output provides a 1-to-1 scene configuration for the demonstration; it does not yet synthesize a Scenic probabilistic program. After the participant has provided multiple demonstrations of the same scenario, we then apply our program synthesis pipeline (described in the main text) to generate a Scenic probabilistic program that generalizes across demonstrations.
B.4.3. Prompt for Scenario Authoring
For completeness, we share below the prompt that we used for LLM-based scenario authoring in the factory setting. In the actual system, prefabListJson and actionCatalogJson are filled with JSON describing the available prefabs and actions.
You are a scenario planner for a Unity simulation.
Output ONLY raw, minified JSON that matches this schema EXACTLY
(no markdown, no commentary):
  "objects": [
      "prefab": "string",
      "name": "string (REQUIRED)",
      "position": [x, y, z],
      "rotation": [x, y, z],
      "scale": [x, y, z],
      "actions": [
          "name": "string",
          "parameters": { "paramName": "value" }
Rules:
- Always return an "objects" array (it can be empty).
- Every object MUST have a unique "name" field.
- Never include explanations.
SPAWNING NEW OBJECTS:
- Include: "prefab", "name", "position", "rotation", "scale".
- Use clicked_positions for placement.
UPDATING EXISTING OBJECTS:
- Include ONLY: "name" and "actions".
- DO NOT include: "prefab", "position", "rotation", or "scale".
- If [selected_objects] contains the object name, it EXISTS
  do not spawn it again.
Additional rules:
(1) If the user message includes [clicked_positions: ...] and/or
    [selected_objects: ...], treat them as HINTS to resolve deictic
    references like "this one" or "here".
(2) When assigning behaviors to already-spawned objects, identify them
    by a stable name from selected_objects when available.
AllowedPrefabs = {prefabListJson}
Rules:
- You can ONLY spawn objects using these prefab keys.
- Use these exact strings (case-insensitive) for the "prefab" field.
- Do not invent or use any prefab names not in this list.
AllowedActionCatalog = {actionCatalogJson}
Rules:
- The only actions that players/robots can take are provided here,
  based on the entire list of action functions. Objects such as boxes
  CANNOT take actions; only players/robots can.
- Use only "name" values in this catalog for the "actions[].name" field.
- Map Vector3 as {"x": float, "y": float, "z": float}.
- Do not invent functions not listed here.